OpenStudy (anonymous):

Algebra 2. Multiple choice? Please help :/

OpenStudy (anonymous):

With the restriction of \[x \neq0\] which of the following is the simplified form of \[\frac{ 6x^4 }{ 4x^9 }\times \frac{ 12x^2 }{ 3x^5 }\]

OpenStudy (anonymous):

\[A. \frac{ 6 }{ x^2 }\]

OpenStudy (anonymous):

\[B. 6x^8\]

OpenStudy (anonymous):

\[C. \frac{ 6 }{ x^8 }\]

OpenStudy (anonymous):

\[D. 6x^2\]

OpenStudy (stamp):

\[\frac{6x^4*12x^2}{4x^9*3x^5}\]

OpenStudy (stamp):

\[\frac{6*12*x^2*x^4}{3*4*x^5*x^9}\]

OpenStudy (stamp):

additional help - http://tutorial.math.lamar.edu/Classes/Alg/IntegerExponents.aspx

OpenStudy (anonymous):

Thanks! I think I got it from here :D

OpenStudy (stamp):

\[\frac{2*3}{x^3*x^5}\]

OpenStudy (stamp):

That should be enough help